In the United States, voters who are neither Democrat nor Republican are called independent. It s survey asked 37 people to identity themselves as Democrat, Republican, or Independent believed that 8% of voters are independent A A What is the probability that none ot the people are Independent? B. What is the probability that tewer than 6 are Independent? Probability C What is the probability that more than 2 people are Independent? Probability
Answer #1

Let X be a number of independent people.

Here, X ~ Binomial ( n = 37, p = 0.08)

Using Excel we get, following probabilities,

A) P(X = 0) = BINOMDIST( 0, 37, 0.08, 0) = 0.0457

Probabilty = 0.0457

B) P(X < 6)=P(X <= 5)=BINOMDIST(5, 37, 0.08, 1)= 0.9285

Probability = 0.9285

C)

P(X > 2)

= 1 - P(X <= 2)

= 1 - BINOMDIST(2, 37, 0.08, 1)

= 1 - 0.4231

= 0.5769

Probability = 0.5769